Solve 195=200-5(1.2)^3x-2
Answer by fractalier(6550)

195 = 200 - 5(1.2)^3x-2
-5 = -5(1.2)^3x-2
1 = (1.2)^3x-2
ln 1 = ln (1.2)^3x-2 = (3x-2)*ln(1.2)
0 = (3x-2)*ln(1.2)
Since a log (or ln) cannot equal zero, 3x - 2 must equal zero, so that
x = 2/3